OpenStudy (anonymous):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)? (x + 2)^2 + (y + 3)^2 = 9 (x − 2)^2 + (y − 3)^2 = 16 (x − 2)^2 + (y − 3)^2 = 4 (x + 2)^2 + (y + 3)^2 = 16 (x − 2)2 + (y + 3)2 = 9

undeadknight26 (undeadknight26):

Hewo der rainbow! Which do you think it is?

OpenStudy (anonymous):

I honestly have no clue what so ever, but if I had to guess I would choose the last one.

OpenStudy (misty1212):

JI!!

undeadknight26 (undeadknight26):

Close...But not right...

OpenStudy (misty1212):

no not the last one don't pick that one

undeadknight26 (undeadknight26):

Last one is horrible choice ^

OpenStudy (anonymous):

Lol, i told you, I have no clue

undeadknight26 (undeadknight26):

Probably @misty1212 can tell you how to do it...I just graph it :p

OpenStudy (misty1212):

\[(x-h)^2+(y-k)^2=r^2\] and in your case \((h,k)\) is \((-2,-3)\) so \[(x+2)^2+(y+3)^2=r^2\]

OpenStudy (anonymous):

Okay, so how would I find the r^2?

OpenStudy (misty1212):

\(r\) is the distance between \((-2,0)\) and \((-2,3)\) which is evidently \(3\) so \[(x+2)^2+(y+3)^2=3^2\]

OpenStudy (anonymous):

Ohhh I get it, so the answer is \[(x+2)^2 + (y+3)^2 = 9 \]

undeadknight26 (undeadknight26):

Yup.

OpenStudy (anonymous):

Ohh, okay thanks to both. That really helped me :)